# Causal Clustering for Conditional Average Treatment Effects Estimation and Subgroup Discovery

### v1

@@ I-A Heterogeneity in Treatment Effects
Causal inference aims to determine the effect of a specific intervention or treatment on an outcome of interest (rubin1974estimating, ; holland1986statistics, ). This is particularly valuable in fields such as healthcare, economics, and social sciences, where accurately estimating the effects of policies or interventions is essential. Using the potential outcomes framework (holland1986statistics, ), we observe a dataset of $n$ independent and identically distributed samples $\{(X_{i},Y_{i},W_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represents individual features, $Y_{i}\in\mathbb{R}$ is the observed outcome, and $W_{i}\in\{0,1\}$ denotes the treatment assignment. For this binary treatment, the average treatment effect (ATE) is defined as $\mathbb{E}[Y^{(1)}-Y^{(0)}]$, where $Y^{(w)}$ represents the potential outcome that would have been observed under treatment $W=w$. However, treatment effects often vary across subgroups in both magnitude and direction. A treatment beneficial for one subgroup may be harmful for another, and reporting only the ATE can obscure such heterogeneity (kravitz2004evidence, ).
@@
A wide range of methods has been developed to estimate CATE, including recursive partitioning (athey2016recursive, ; athey2019generalized, ; friedberg2020local, ), partially linear models (robins1992estimating, ), and neural networks (shalit2017estimating, ), among others. Additionally, meta-learners, such as the X-Learner (kunzel2019metalearners, ) and the R-Learner (nie2021quasi, ), integrate multiple approaches to enhance estimation accuracy.
@@ I-B Motivation and Current Work
While clustering is a well-established tool in unsupervised learning, its application to causal inference remains relatively nascent. The key challenge stems from the fundamental problem of causal inference (holland1986statistics, ): we never observe both treated and untreated outcomes for the same individual, necessitating an estimation step before any subgrouping can be applied. This leads to a two-stage procedure: first estimating individual treatment effects, then analyzing their structure to uncover latent subpopulations with heterogeneous responses.
Several strands of related work have pursued this two-stage strategy. One direction focuses on clustering with latent or noisy data, such as Gaussian mixture models trained via the Expectation-Maximization (EM) algorithm (dempster1977maximum, ; serafini2020handlingmissingdatamodelbased, ; sportisse2021model, ; boluki2019optimal, ), and model-based clustering under measurement error (kumar2007clustering, ; zhang2020model, ). These lines of work, while not restricted to causal inference in application, do not handle adjusting estimates when propensity is unbalanced. More recently, (lee2024subgroupteadvancingtreatmenteffect, ) proposed a joint EM framework that clusters based on radial basis function (RBF) distances between CATE estimates, with the distance weights learned via a transformer architecture. Another direction involves explicitly estimating the CATE and then assigning subgroups through decision rules or clustering. For example, (foster2011subgroup, ; bargaglistoffi2024causalruleensembleinterpretable, ; argaw2022identifyingheterogeneoustreatmenteffects, ) perform recursive partitioning on the estimated CATE surface to derive interpretable subgroup assignments. For the tree based rules such as CRE by (bargaglistoffi2024causalruleensembleinterpretable, ), there is no direct way to set the number of desired clusters. In our work, we believe in allowing users to flexibly incorporate a variety of existing libraries and solvers at each step.
Interestingly, while tree-based methods such as classification and regression trees, random forests (breiman2001random, ), and causal forests (athey2019generalized, ) are widely used for both CATE estimation and subgroup discovery, often by leveraging splitting rules, there has been limited exploration of using these ensembles as explicit kernel density estimators for subgroup discovery, which presents an opportunity to incorporate kernel learning techniques. Forest-based methods are well known to induce similarity metrics that behave like adaptive local kernel estimators (hothorn2004bagging, ; fugon2008data, ; athey2019generalized, ; friedberg2020local, ). In the traditional unsupervised learning literature, these forest-induced kernels have even been used in spectral clustering algorithms (Yan_2013, ), suggesting potential synergy in the causal setting that motivates this research.
@@ I-C Our Contribution: A New Perspective
Therefore, in this work, we introduce a novel perspective on subgroup discovery for heterogeneous treatment effects that informs the clustering process using a learned kernel (similarity matrix) derived from residual-on-residual (R-Learner) style estimators (nie2021quasi, ). We frame the hard clustering step as a form of implicit regularization on the CATE estimation task, promoting smoothness and subgroup coherence in the estimated treatment effects. We view this work as a first step toward a broader class of tools that combine kernel learning and causal effect estimation in a modular and interpretable way. We believe this method represents a promising direction for exploratory analysis of treatment effect heterogeneity.
@@ II-A The Robinson Decomposition of the CATE
In order to identify the CATE $\tau(x)$, we rely on the following commonly used identification strategies in (imbens2015causal, ):
@@
We now review the Robinson decompostion by (robinson1988root, ). To aid readability, we rewrite the following terms below as:
@@
This decomposition, originally used by (robinson1988root, ) for partially linear models, has seen a considerable revival of interest in recent years and it has been used in many popular and seminal works such as the Double Machine Learning (DML) framework (chernozhukov2018double, ) and forest ensemble methods such as the Orthogonal Random Forest (ORF) (oprescu2019orthogonal, ), Causal Forests (CF) (athey2019generalized, ). In particular, 6 can be re-formulated as a regularized empirical loss minimization program below in Equation (7):
@@
Here $\Lambda$ is a regularization term that penalizes the complexity of the functional form of the CATE decision variable $\tau$ either explicitly (e.g., penalized regression) or implicitly (e.g., complexity of a neural network, tree depths of a forest). This residual-on-residual regression approach has incredible flexibility in allowing the propensity $e(\cdot)$ and mean outcomes $m(\cdot)$ to be estimated separately with a variety of supervised machine learning techniques. In addition, it has very attractive statistical properties when the estimation is done with proper sample-splitting and cross-fitting techniques (cf. DML by (chernozhukov2018double, ), Neyman Orthogonalization by (foster2023orthogonal, )). For our work, we are interested in building upon the R-Learner framework introduced by (nie2021quasi, ) through the causal forest first introduced by (athey2016recursive, ). The former specifically studies the properties for Equation (7) when $\tau$ belongs to the Reproducing Kernel Hilbert Space (RKHS) induced by a positive semi-definite kernel function $K$ and thus admits a large variety of underlying estimators such as linear models, kernel ridge regression, and even more complex non/semi-parametric models such as ensembles. Tree ensemble methods frame the estimation problem as a locally weighted regression task, where the weights are implicitly defined by the frequency with which a sample falls into the same leaf as other samples (referred to as forest weights). These forest ensemble methods minimize the residualized loss around an input sample with features $x$ as follows:
@@
where $\alpha_{j}(x)$ represents the forest weights matrix (which will be formally defined in Section III). Notably, both the response function $\hat{m}_{x}$ and the propensity function $\hat{e}_{x}$ are locally estimated, as indicated by the subscripts with $x$, and debiased by means of cross-fitting (chernozhukov2018double, ) (also called “honest-fitting”). The regularization penalty term $\Lambda(\tau(x))$ controls the complexity of the treatment effect function $\tau(\cdot)$, and it is implicitly determined by various tree parameters such as maximum depth, minimum leaf size, and tree balance. Crucially, this framework enables the learning of a kernel that captures the similarity between sample points based on their Conditional Average Treatment Effect (CATE). This kernel provides an intuitive mapping of the relationship between samples, which is instrumental in the subsequent step of kernelized clustering.
@@ II-B Clustered Orthogonal Learner
While the above description is typically more intuitive, it is more convenient and insightful to formulate our clustering problem via the equivalent Sum-of-Norms (SON) model (pelckmans2005convex, ) as follows:
@@
where the penalty term, controlled by $\lambda\geq 0$, enforces similarity between treatment effect estimates across samples. The choice of $q=0$ (pseudo-norm) explicitly limits the number of unique centroids, effectively determining the number of clusters. When $\lambda=0$, the model reduces to the original un-clustered orthogonal learner, while increasing $\lambda$ encourages greater clustering, reducing the number of distinct centroids (with $\lambda\rightarrow\infty$ forcing the estimator $\hat{\tau}$ to output the same estimate for every $X_{i}$). This reformulation also aligns itself with the residual-on-residual structure of the R-Loss in Equation (7), with the number of clusters i.e., total number of unique CATE estimates forming a natural regularizing component. Nevertheless, it presents a computational and theoretical challenge as optimal k-means clustering—an instance of the Minimum Sum-of-Squares Clustering (MSSC) problem—is known to be NP-hard (aloise2009np, ). To address this, we introduce a relaxation of Equation (9) along with a practical implementation in the following section.
@@ III Proposed Procedure
Computationally, this sequential approach relaxes the original problem in two key ways: First it transforms the simultaneous optimization of treatment effect estimation and clustering into a more tractable sequential process. Second, the NP-hard, non-convex clustering problem is replaced with clustering methods that can be solved using traditional iterative procedures e.g., Lloyd’s algorithm (MacQueen1967, ) or as a semi-definite program via methods by (chi2015splitting, ; sun2021convex, ; touw2022convex, ).
By structuring the optimization in this manner, we balance computational feasibility with the flexibility of kernel-based clustering while preserving the algorithmic flexibility of the initial estimation step, making implementation with out-of-the box libraries straightforward. More importantly, by decoupling the initial learning step and subsequent clustering step and allowing for honest estimation, the clustering analysis (Step 2) can also be conducted on a new set of unseen test data points without retraining the estimators of (Step 1), resulting in reduced inference times. This modification also allows us to conduct honest estimation (chernozhukov2018double, ; nie2021quasi, ; athey2019generalized, ; foster2023orthogonal, ), where separate data subsets are used in cluster construction and estimation, thus avoiding overfitting and preserving the integrity of subgroup discovery. We now describe how to obtain a valid learned kernel from the trained $\tau(\cdot)$ estimator.
@@ III-A Forest Based Kernels
To effectively utilize the capabilities of tree-based ensemble methods, we leverage the grf package in R (athey2019generalized, ). Specifically, we can construct a valid kernel matrix by extracting and transforming the “forest weights matrix” from a trained forest model. Given a forest of $B$ trees trained on $n$ data points and $m$ new test data points, which may include previously unseen instances, we derive a non-negative $m\times n$ similarity matrix $\alpha$ that quantifies the relationship between each new test data point with each of the training data points. This relationship is mathematically expressed as follows, adhering to the notation established by (athey2019generalized, ) and (friedberg2020local, ) in the following form:
@@ IV-A Methods
We evaluate the performance of our proposed clustering methodology on both synthetic and real-world datasets. First, in §IV-B, we conduct an ablation study using a semi-synthetic variant of the widely studied Infant Health and Development Program (IHDP) dataset (DVN/D9TBWP_2017, ), assessing how different kernel constructions and cross-fitting strategies affect CATE estimation and subgroup quality. Next, in §IV-C, we tackle the problem of cluster size selection by evaluating the performance of various cluster size selection methods on a synthetic dataset. Following that, §IV-D investigates a synthetic adversarial design in which the data is deliberately constructed without any true underlying clusters, to investigate performance under model misspecification with complex response surfaces and high noise. Finally, in §IV-E, we apply our method to the Synthea (walonoski2018synthea, ; lee2025accurate, ) dataset involving emergency health records with confounded treatment assignments. This is to demonstrate its ability to provide relevant insights in real-world scenarios beyond randomized controlled trials.
@@ IV-B Semi-synthetic Ablation Study
Our ablation study aims to evaluate the impact of kernel construction and estimation strategies on the quality of CATE-based clustering. The IHDP dataset originates from a randomized clinical trial aimed at improving the cognitive and health outcomes of low-birth-weight, premature infants through intensive child care and home visits. The dataset includes 747 individuals, of whom 139 received the treatment and 608 did not, each characterized by 25 covariates capturing demographic, clinical, and socio-economic factors. The version we use is a semi-synthetic dataset constructed by retaining the original covariates from the study while simulating potential outcomes according to predefined response surfaces and is widely used in causal inference research (cf. (lee2020robustrecursivepartitioningheterogeneous, ; argaw2022identifyingheterogeneoustreatmenteffects, ; DBLP:journals/corr/AlaaS17, ; pmlr-v70-shalit17a, )).
Our default method is Cross Fitted, which uses out-of-fold predictions to construct the kernel matrix from learned CATE weights, mitigating overfitting and improving generalization. We compare this against several ablations. First, the Thresholded method mimics the “thresholding” by CATE type of heuristic approach, where the RBF matrix of features is binarized at the 90th percentile (10.5555/2986459.2986566, ). Next, the RBF Cross Fitted variant applies spectral clustering on RBF kernels of raw features instead of the cross fitted forest kernel. Then, the RBF Oracle method uses ground-truth CATE values to compute the idealized cluster-level statistics on the RBF of features to demonstrate the importance of clustering on the learnt kernel instead. Finally, we also evaluate the CRE package by (bargaglistoffi2024causalruleensembleinterpretable, ), a recursive partitioning based method.
@@ IV-C Cluster Size Selection and Recovery Simulation
For each run above, we apply our framework to first estimate, then cluster on the learnt kernel. We try out 4 popular methods for cluster size selection: The eigengap (ng2001spectral, ), elbow (thorndike1953belongs, ), silhouette (rousseeuw1987silhouettes, ), and the gap statistic (gap_stat) (tibshirani2001estimating, ). We evaluate the quality of the cluster recovery results via the Adjusted Rand Index (ARI) and the Normalized Mutual Information (NMI) (see Appendix Appendix A: Clustering Metrics Definitions for definition). We present the results of the cluster recovery simulation design in Table II and display one example of a cluster recovery result in Figure 1 for illustration purposes.

### v2

@@ I-A Heterogeneity in Treatment Effects
Causal inference aims to determine the effect of a specific intervention or treatment on an outcome of interest [1, 2]. This is particularly valuable in fields such as healthcare, economics, and social sciences, where accurately estimating the effects of policies or interventions is essential. Using the potential outcomes framework [2], we observe a dataset of $n$ independent and identically distributed samples $\{(X_{i},Y_{i},W_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represents individual features, $Y_{i}\in\mathbb{R}$ is the observed outcome, and $W_{i}\in\{0,1\}$ denotes the treatment assignment. For this binary treatment, the average treatment effect (ATE) is defined as $\mathbb{E}[Y^{(1)}-Y^{(0)}]$, where $Y^{(w)}$ represents the potential outcome that would have been observed under treatment $W=w$. However, treatment effects often vary across subgroups in both magnitude and direction. A treatment beneficial for one subgroup may be harmful for another, and reporting only the ATE can obscure such heterogeneity [3].
@@
A wide range of methods has been developed to estimate CATE, including recursive partitioning [4, 5, 6], partially linear models [7], and neural networks [8], among others. Additionally, meta-learners, such as the X-Learner [9] and the R-Learner [10], integrate multiple approaches to enhance estimation accuracy.
@@ I-B Motivation and Current Work
While clustering is a well-established tool in unsupervised learning, its application to causal inference remains relatively nascent. The key challenge stems from the fundamental problem of causal inference [2]: we never observe both treated and untreated outcomes for the same individual, necessitating an estimation step before any subgrouping can be applied. This leads to a two-stage procedure: first estimating individual treatment effects, then analyzing their structure to uncover latent subpopulations with heterogeneous responses.
Several strands of related work have pursued this two-stage strategy. One direction focuses on clustering with latent or noisy data, such as Gaussian mixture models trained via the Expectation-Maximization (EM) algorithm [11, 12, 13, 14], and model-based clustering under measurement error [15, 16]. These lines of work, while not restricted to causal inference in application, do not handle adjusting estimates when propensity is unbalanced. More recently, [17] proposed a joint EM framework that clusters based on radial basis function (RBF) distances between CATE estimates, with the distance weights learned via a transformer architecture. Another direction involves explicitly estimating the CATE and then assigning subgroups through decision rules or clustering. For example, [18, 19, 20] perform recursive partitioning on the estimated CATE surface to derive interpretable subgroup assignments. For the tree based rules such as CRE by [19], there is no direct way to set the number of desired clusters. In our work, we believe in allowing users to flexibly incorporate a variety of existing libraries and solvers at each step.
Interestingly, while tree-based methods such as classification and regression trees, random forests [21], and causal forests [5] are widely used for both CATE estimation and subgroup discovery, often by leveraging splitting rules, there has been limited exploration of using these ensembles as explicit kernel density estimators for subgroup discovery, which presents an opportunity to incorporate kernel learning techniques. Forest-based methods are well known to induce similarity metrics that behave like adaptive local kernel estimators [22, 23, 5, 6]. In the traditional unsupervised learning literature, these forest-induced kernels have even been used in spectral clustering algorithms [24], suggesting potential synergy in the causal setting that motivates this research.
@@ I-C Our Contribution: A New Perspective
Therefore, in this work, we introduce a novel perspective on subgroup discovery for heterogeneous treatment effects that informs the clustering process using a learned kernel (similarity matrix) derived from residual-on-residual (R-Learner) style estimators [10]. We frame the hard clustering step as a form of implicit regularization on the CATE estimation task, promoting smoothness and subgroup coherence in the estimated treatment effects. We view this work as a first step toward a broader class of tools that combine kernel learning and causal effect estimation in a modular and interpretable way. We believe this method represents a promising direction for exploratory analysis of treatment effect heterogeneity.
@@ II-A The Robinson Decomposition of the CATE
In order to identify the CATE $\tau(x)$, we rely on the following commonly used identification strategies in [25]:
@@
We now review the Robinson decompostion by [26]. To aid readability, we rewrite the following terms below as:
@@
This decomposition, originally used by [26] for partially linear models, has seen a considerable revival of interest in recent years and it has been used in many popular and seminal works such as the Double Machine Learning (DML) framework [27] and forest ensemble methods such as the Orthogonal Random Forest (ORF) [28], Causal Forests (CF) [5]. In particular, 6 can be re-formulated as a regularized empirical loss minimization program below in Equation (7):
@@
Here $\Lambda$ is a regularization term that penalizes the complexity of the functional form of the CATE decision variable $\tau$ either explicitly (e.g., penalized regression) or implicitly (e.g., complexity of a neural network, tree depths of a forest). This residual-on-residual regression approach has incredible flexibility in allowing the propensity $e(\cdot)$ and mean outcomes $m(\cdot)$ to be estimated separately with a variety of supervised machine learning techniques. In addition, it has very attractive statistical properties when the estimation is done with proper sample-splitting and cross-fitting techniques (cf. DML by [27], Neyman Orthogonalization by [29]). For our work, we are interested in building upon the R-Learner framework introduced by [10] through the causal forest first introduced by [4]. The former specifically studies the properties for Equation (7) when $\tau$ belongs to the Reproducing Kernel Hilbert Space (RKHS) induced by a positive semi-definite kernel function $K$ and thus admits a large variety of underlying estimators such as linear models, kernel ridge regression, and even more complex non/semi-parametric models such as ensembles. Tree ensemble methods frame the estimation problem as a locally weighted regression task, where the weights are implicitly defined by the frequency with which a sample falls into the same leaf as other samples (referred to as forest weights). These forest ensemble methods minimize the residualized loss around an input sample with features $x$ as follows:
@@
where $\alpha_{j}(x)$ represents the forest weights matrix (which will be formally defined in Section III). Notably, both the response function $\hat{m}_{x}$ and the propensity function $\hat{e}_{x}$ are locally estimated, as indicated by the subscripts with $x$, and debiased by means of cross-fitting [27] (also called “honest-fitting”). The regularization penalty term $\Lambda(\tau(x))$ controls the complexity of the treatment effect function $\tau(\cdot)$, and it is implicitly determined by various tree parameters such as maximum depth, minimum leaf size, and tree balance. Crucially, this framework enables the learning of a kernel that captures the similarity between sample points based on their Conditional Average Treatment Effect (CATE). This kernel provides an intuitive mapping of the relationship between samples, which is instrumental in the subsequent step of kernelized clustering.
@@ II-B Clustered Orthogonal Learner
While the above description is typically more intuitive, it is more convenient and insightful to formulate our clustering problem via the equivalent Sum-of-Norms (SON) model [30] as follows:
@@
where the penalty term, controlled by $\lambda\geq 0$, enforces similarity between treatment effect estimates across samples. The choice of $q=0$ (pseudo-norm) explicitly limits the number of unique centroids, effectively determining the number of clusters. When $\lambda=0$, the model reduces to the original un-clustered orthogonal learner, while increasing $\lambda$ encourages greater clustering, reducing the number of distinct centroids (with $\lambda\rightarrow\infty$ forcing the estimator $\hat{\tau}$ to output the same estimate for every $X_{i}$). This reformulation also aligns itself with the residual-on-residual structure of the R-Loss in Equation (7), with the number of clusters i.e., total number of unique CATE estimates forming a natural regularizing component. Nevertheless, it presents a computational and theoretical challenge as optimal k-means clustering—an instance of the Minimum Sum-of-Squares Clustering (MSSC) problem—is known to be NP-hard [31]. To address this, we introduce a relaxation of Equation (9) along with a practical implementation in the following section.
@@ III Proposed Procedure
Computationally, this sequential approach relaxes the original problem in two key ways: First it transforms the simultaneous optimization of treatment effect estimation and clustering into a more tractable sequential process. Second, the NP-hard, non-convex clustering problem is replaced with clustering methods that can be solved using traditional iterative procedures e.g., Lloyd’s algorithm [32] or as a semi-definite program via methods by [33, 34, 35].
By structuring the optimization in this manner, we balance computational feasibility with the flexibility of kernel-based clustering while preserving the algorithmic flexibility of the initial estimation step, making implementation with out-of-the box libraries straightforward. More importantly, by decoupling the initial learning step and subsequent clustering step and allowing for honest estimation, the clustering analysis (Step 2) can also be conducted on a new set of unseen test data points without retraining the estimators of (Step 1), resulting in reduced inference times. This modification also allows us to conduct honest estimation [27, 10, 5, 29], where separate data subsets are used in cluster construction and estimation, thus avoiding overfitting and preserving the integrity of subgroup discovery. We now describe how to obtain a valid learned kernel from the trained $\tau(\cdot)$ estimator.
@@ III-A Forest Based Kernels
To effectively utilize the capabilities of tree-based ensemble methods, we leverage the grf package in R [5]. Specifically, we can construct a valid kernel matrix by extracting and transforming the “forest weights matrix” from a trained forest model. Given a forest of $B$ trees trained on $n$ data points and $m$ new test data points, which may include previously unseen instances, we derive a non-negative $m\times n$ similarity matrix $\alpha$ that quantifies the relationship between each new test data point with each of the training data points. This relationship is mathematically expressed as follows, adhering to the notation established by [5] and [6] in the following form:
@@ IV-A Methods
We evaluate the performance of our proposed clustering methodology on both synthetic and real-world datasets. First, in §IV-B, we conduct an ablation study using a semi-synthetic variant of the widely studied Infant Health and Development Program (IHDP) dataset [36], assessing how different kernel constructions and cross-fitting strategies affect CATE estimation and subgroup quality. Next, in §IV-C, we tackle the problem of cluster size selection by evaluating the performance of various cluster size selection methods on a synthetic dataset. Following that, §IV-D investigates a synthetic adversarial design in which the data is deliberately constructed without any true underlying clusters, to investigate performance under model misspecification with complex response surfaces and high noise. Finally, in §IV-E, we apply our method to the Synthea [37, 38] dataset involving emergency health records with confounded treatment assignments. This is to demonstrate its ability to provide relevant insights in real-world scenarios beyond randomized controlled trials.
@@ IV-B Semi-synthetic Ablation Study
Our ablation study aims to evaluate the impact of kernel construction and estimation strategies on the quality of CATE-based clustering. The IHDP dataset originates from a randomized clinical trial aimed at improving the cognitive and health outcomes of low-birth-weight, premature infants through intensive child care and home visits. The dataset includes 747 individuals, of whom 139 received the treatment and 608 did not, each characterized by 25 covariates capturing demographic, clinical, and socio-economic factors. The version we use is a semi-synthetic dataset constructed by retaining the original covariates from the study while simulating potential outcomes according to predefined response surfaces and is widely used in causal inference research (cf. [39, 20, 40, 41]).
Our default method is Cross Fitted, which uses out-of-fold predictions to construct the kernel matrix from learned CATE weights, mitigating overfitting and improving generalization. We compare this against several ablations. First, the Thresholded method mimics the “thresholding” by CATE type of heuristic approach, where the RBF matrix of features is binarized at the 90th percentile [42]. Next, the RBF Cross Fitted variant applies spectral clustering on RBF kernels of raw features instead of the cross fitted forest kernel. Then, the RBF Oracle method uses ground-truth CATE values to compute the idealized cluster-level statistics on the RBF of features to demonstrate the importance of clustering on the learnt kernel instead. Finally, we also evaluate the CRE package by [19], a recursive partitioning based method.
@@ IV-C Cluster Size Selection and Recovery Simulation
For each run above, we apply our framework to first estimate, then cluster on the learnt kernel. We try out 4 popular methods for cluster size selection: The eigengap [43], elbow [44], silhouette [45], and the gap statistic (gap_stat) [46]. We evaluate the quality of the cluster recovery results via the Adjusted Rand Index (ARI) and the Normalized Mutual Information (NMI) (see Appendix Appendix A: Clustering Metrics Definitions for definition). We present the results of the cluster recovery simulation design in Table II and display one example of a cluster recovery result in Figure 1 for illustration purposes.